The vapour density of mixture containig NO2 and N2O4 is 27.6.The mole fraction of N2O4 is?

Grade:11

3 Answers

Arun
6 years ago
Dear student

Molar mass of the mixture = 27.6 x 2 = 55.2 g
Let x moles of N2o4 be present in the 100 mole mixture
So,(100-x) x 46 + 92x = 100 x 55.2
=> 100 - x + 2x = 5520/46 = 120
=> 100 + x = 120
=> x = 20
Hence,mole fraction of N2O4 in the mixture = 20/100 = 0.2
Regards
Arun (askIITians forum expert)
Vishal Kumar
2 years ago
Degree of dissociation = D/d -1 = 46/27.6 -1 = 18.4/27.6
[N2O4] = 1- [18.4/27.6] = 1/3
Total mole = 1+ [18.4/27.6] = 46 / 27.6
Mole fraction of N2O4 = 1/3 × 27.6/46 = 0.2
Ujjawal Mishra
2 years ago
 All  are telling anwer in very tough way. We cannot understand even a single step. Its waste of time for us.